Percent Yield question:

Glucose into ethanol and carbon dioxide is given:

C6H12O6 -> C2H5OH + CO2


If 24g of ethanol (c2h5oh) was recovered from the conversion of 50g of glucose (c6h12o6)..


1.) What is the percent yield?


1
Expert's answer
2021-04-25T07:47:00-0400

The balanced reaction equation:

C6H12O6 = 2 C2H5OH + 2 CO2

Molar weight of glucose: 180 g/mol

Quantity of glucose: 50 g / (180 g/mol) = 0.278 mol

Theoretical quantity of ethanol: 2 * 0.278 mol = 0.556 mol

Molar weight of ethanol: 46 g/mol

Theoretical weight of ethanol: 0.556 mol * (46 g/mol) = 25.58 g

Yield: (24 g / 25.58 g) * 100% = 93.8%

Answer: 93.8%